Grade 12th passAlgebra

Using appropriate properties find.
-2/3 × 3/5 + 5/2 – 3/5 × 1/6

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

-2/3 × 3/5 + 5/2–3/5 × 1/6
= -2/3 × 3/5–3/5 × 1/6+ 5/2 (by commutativity)
= 3/5 (-2/3–1/6)+ 5/2
= 3/5 ((- 4–1)/6)+ 5/2
= 3/5 ((–5)/6)+ 5/2 (by distributivity)
=-15/30+5/2
=-1/2+5/2
= 4/2
=2

Thanks